OpenStudy (anonymous):

juanita and lauran are painting a circular mural on one wall of the community center. the area of the mural is 33 square meters. what is the radius? use pi=22/7

OpenStudy (anonymous):

The area of a circle is equal to pi times the square of the radius: \[A=\pi r^2\] Plug in our values and solve for r: \[33 m^2 = (22/7)r^2\] \[r=\sqrt{(7/22)(33)}=\sqrt{10.5} \approx 3.24m\]

OpenStudy (anonymous):

my choices are 7/2 or 4 or \[\sqrt{4}\]

OpenStudy (anonymous):

None of those answers are correct, but the closest of those choices is 7/2.

OpenStudy (anonymous):

thanks